N. What is q2 (magnitude and sign)? them. Each object
Answer #1

F = k q1q2/ r^2

2.6 = 9*10^9*3.6**10^-6* q2 / 0.3^2

q2 = 7.22 *10^-6 C

since the charges attracts each other so it must be negative.

======

F = k q^2 / r^2

F = k ( ne) ^2 / r^2

F = k n^2 e^2 / r^2

4.55*10^-21 = 9*10^9* n^2 * ( 1.6*10^-19)^2 / ( 5.4*10^-3)^2

n = 24

========

Comment in case any doubt, will reply for sure. Goodluck